Shift Summary for Owl shift, 9/Jul/99

Miguel Barrio (
Fri, 9 Jul 1999 18:04:03 -0500

Shift Summary for Owl shift, 09/Jul/99

On Shift: Julie Whitmore; Miguel Barrio

Hours of beam available at all intensities: 2.0
Hours of beam available at nominal intensity: 1.0
Hours of E832 data (special runs not included): 1.3
Hours of special runs (laser scans, muon runs...): 0.1

Accounting of lost time: Beam down from 1:20 to 6:30

Runs ended this shift:
Run 13838 - 20.8M events - BEAM_832
Run 13839 - Laser Scan
Run 13840 - VERY short BEAM_832 run, no good
Run 13841 - 120K events - BEAM_MU_4E12

Tapes brought to FCC this shift: None
(including tapes for online split)

Accesses during shift: None

Overview of shift:
Some E832 datataking
Beam went down at around 1:20am
Took laser scan (2:00am-2:45am)

Description of Shift:

run spill time comment
13838 574 ~1:20 Regenerator isn't moving, we call MCR to inquire
(since the reset procedure gives an error).
TeV quench righ after call.

578 We get the regenerator to move again.

597 1:52 TeV message: TeV down, injection quench,
est 30 minutes for beam. We stop the run to
start a laser scan.

13839 2:00 Laser scan run starts.
2:45 End of laser scan.
Beam should be up by 3am (or so the screen

13840 2:58 Start run 13840. No beam yet.
3:03 Message oon screen: "MI LLRF problems, est.
45 min." So no beam for a while.
3:51 Message: "MI LLRF problems, est. 1 hr."
4:12 Got beam back. Died after a couple of minutes.
4:15 Got a vacuum problem. Doug Jensen was paged.
The blower pump shut itself off for some reason,
Doug turned the other pump back on, pump #1
wouldn't come back on, same as diffusion pump #1
6:25 Turned magnet down for muon run. Will turn
magnet back up with the opposite polarity after
the muon run.
6:30 Beam came back (muon mode).

13841 6:40 Start of muon run.
6:45 End of muon run.

13842 6:55 Start of hadron run (BEAM_832).

Follow up Notes on Previous Shift

Items to Follow up on Next Shift